OpenStudy (sh3lsh):

If the arc length of a circle is 17.4 com and the radius of a circle is 5.8 cm, find the area of the circle.

OpenStudy (anonymous):

radius = 5.8cm therefore area of circle = \[\pi r ^{2}\] =22/7 *5.8 *5.8 =105.725 cm square =105.73 cm square(approx) !!! I think that u meant to ask area of the sector and not the area of the circle but still this is the area of the circle !!!

OpenStudy (sh3lsh):

hahah. yeah I thought that, guess the question is messed up. Still, thanks!